# DRPruning: Efficient Large Language Model Pruning through Distributionally Robust Optimization

### v1

@@ Abstract
Large language models (LLMs) deliver impressive results but face challenges from increasing model sizes and computational costs. Structured pruning reduces model size and speeds up inference but often causes uneven degradation across domains, leading to biased performance. To address this, we propose DRPruning, which incorporates distributionally robust optimization to restore balanced performance across domains, along with further improvements to enhance robustness. Experiments in monolingual and multilingual settings show that our method surpasses similarly sized models in pruning and continued pretraining over perplexity, downstream tasks, and instruction tuning. We further provide analysis demonstrating the robustness of our method towards various domains and distribution shifts. Furthermore, our method automatically determines optimal reference losses and data ratios, suggesting potential for broader applications. Our code is available at https://github.com/hexuandeng/DRPruning.
@@
Hexuan Deng1    Wenxiang Jiao    Xuebo Liu1thanks:   Corresponding Author    Min Zhang1    Zhaopeng Tu 1 Institute of Computing and Intelligence, Harbin Institute of Technology, Shenzhen, China {hxuandeng,wenxiangjiaonju,tuzhaopeng}@gmail.com, {liuxuebo,zhangmin2021}@hit.edu.cn
@@ 1 Introduction
Large language models (LLMs) have advanced rapidly, achieving impressive results across a wide range of tasks (Bang et al., 2023; Jiao et al., 2023; Frieder et al., 2023; Bian et al., 2024). However, this progress has come with increasing model sizes, significantly raising computational costs for both training and inference, which impacts their accessibility. Structured pruning is a promising approach to reduce model size (Han et al., 2015; Wen et al., 2016), but it often causes uneven performance degradation across domains, leading to biased capabilities and unfair downstream task performance (Xia et al., 2024).
To address this, we propose a distributionally robust pruning method that dynamically adjusts the data distribution during training to ensure balanced performance across domains. Building on distributionally robust optimization (DRO; Oren et al., 2019; Sagawa et al., 2019), which aims to optimize worst-case performance across distributions, a reference loss is defined for each domain as a target. A higher weight is then assigned to domains with larger deviations from this reference loss, while not straying too far from a predefined reference data ratio. However, setting these hyperparameters is challenging, and suboptimal configurations often result in poor outcomes (Zhou et al., 2021).
To resolve this, we introduce a method that dynamically adjusts hyperparameters during training. Domains with worse performance often exhibit higher losses, deviating further from the expected loss after sufficient training. Using scaling laws (Kaplan et al., 2020; Ghorbani et al., 2022), we predict the loss after training as the reference, where larger deviations indicate poorer performance, thereby promoting capability recovery in these areas. Additionally, we gradually increase the reference data ratio for domains with greater deviations, ensuring robustness across a wider range of distributions, particularly more challenging ones.
We validate our approach through experiments in monolingual and multilingual settings, which represent varying degrees of distributional shift. Our model consistently outperforms Xia et al. (2024) and other LLMs in both pruning and continued pretraining, as measured by perplexity (-5.59% and -5.60%, respectively), downstream tasks (+1.27% and +1.78%, respectively), and instruction tuning (55.4% win rate). Particularly in multilingual settings, our model achieves +2.95% in downstream tasks. To further assess domain-specific performance, we develop a sentence continuation benchmark using existing unlabeled data. Results show strong consistency with downstream task performance, confirming the benchmark’s reliability and further demonstrating our improved domain-level capabilities (+17.9% on average). Moreover, our approach to determining reference losses and data ratios could potentially enhance data efficiency across a broader range of applications, driving further advancements in LLMs.
@@
  • We tackle domain imbalance in structured pruning by introducing a distributionally robust pruning method, that dynamically adjusts data ratios during training to ensure robustness against distributional shifts.
  • We demonstrate the method’s effectiveness through extensive experiments in monolingual and multilingual settings. Our method surpasses existing ones in perplexity, downstream tasks, instruction tuning, and domain-specific evaluation.
  • We offer refined reference losses and data ratios for training, which can be applied more broadly to enhance various model training processes and contribute to advancements for LLMs.
@@ 2.1 Structured Pruning
To prune the model to any target configuration, we adopt structured pruning based on Sheared Llama (Xia et al., 2024). For each granularity $i$, we learn pruning masks $Z=\{\mathbf{z}^{i}\mid\mathbf{z}^{i}\in\mathbb{R}^{D_{i}}\}$ that determine whether substructures are pruned or retained, where $z^{i}_{j}=0$ indicates pruning of the $j$-th substructure. We apply pruning across granularities that include transformer layers, hidden dimensions, attention heads, and FFN intermediate dimensions.
To parameterize the masks, we use the $\ell_{0}$ regularization method (Louizos et al., 2018) with hard concrete distributions that concentrate probability mass at 0 or 1. We further use Lagrange multipliers to ensure the pruned model meets the target configuration. Specifically, if exactly $t^{i}$ parameters must be retained for $\mathbf{z}^{i}$, we impose the constraint:
~$\tilde{\ell}^{i}=\lambda^{i}\left(\sum_{j}z^{i}_{j}-t^{i}\right)+\phi^{i}\left(\sum_{j}z^{i}_{j}-t^{i}\right)^{2}.$ (1)
The final training loss integrates these constraints with the language modeling loss of the pruned model, jointly optimizing the model parameters $\theta$ and pruning masks $\mathbf{z}$, where $\mathbf{z}$ typically uses a higher learning rate. After pruning, we retain the highest-scoring components.
@@ 2.2 Distributionally Robust Optimization
To mitigate uneven domain performance after pruning, we apply distributionally robust optimization (DRO; Oren et al., 2019; Sagawa et al., 2019) to improve the model’s robustness to distribution shifts. DRO seeks a model $\theta$ that performs well across a set of potential test distributions $\mathcal{Q}$ over $n$ domains. Formally, we aim to:
@@
To solve the min-max optimization, we use the iterative best response (IBR) algorithm (Fudenberg and Levine, 1998). Each iteration consists of first performing Empirical Risk Minimization (ERM) on the current data distribution $\mathbf{q}^{t}$, followed by updating the data distribution using worst-case weights based on the current parameters. Formally,
@@
3 Methodology
In Section 3.1, we introduce our overall procedure to utilize these components to design an effective pruning method, as shown in Figure 1. Further, to determine the data ratio $\mathbf{q}^{t}$, DRO requires two inputs: loss function $\ell\left(\theta;D\right)$ and potential distributions $\mathcal{Q}$, which strongly influence $\mathbf{q}^{t}$ and, consequently, the model’s training process. Therefore, we discuss how to better calculate these parameters in Section 3.2 and 3.3.
@@ 3.1 Distributionally Robust Pruning
Building on the aforementioned methods, we construct a pruning approach to enhance the model’s robustness to both capacity reduction and distribution shifts. Specifically, we adopt common techniques (Ma et al., 2023; Li et al., 2024; Zhang et al., 2024b), first applying structured pruning to reduce model parameters, followed by continued pretraining to restore capabilities. Compared to training from scratch, this approach requires only a small amount of unlabeled data to restore the model’s performance (Xia et al., 2024).
Combining DRO and pruning.
We further integrate the DRO method into both the pruning and continued pretraining stages. During training, DRO dynamically adjusts the data ratio to improve the model’s robustness and convergence speed. Specifically, to prevent overfitting, we compile a validation set and use the evaluation loss as the loss score. After each evaluation (every 50 steps during pruning and every 400 steps during continued pretraining), we update the data ratio based on the evaluation loss using the DRO method, directing training to focus more on underperforming domains.
@@
Next, we optimize the loss function (Section 3.2) and potential distributions (Section 3.3) to ensure robust training, as shown in Figure 1. In contrast, Sheared Llama employs a dynamic scheduling strategy that forces the model to strictly adhere to the relative loss magnitudes of larger LLMs, without placing any constraints on the potential distributions. This leads to suboptimal results, particularly in multilingual settings with significant distribution shifts.
@@ 3.2 Dynamic Loss Function
^$\hat{\ell}(P,T)=A\cdot\frac{1}{P^{\alpha}}\cdot\frac{1}{T^{\beta}}+E,$ (4)
@@ 3.3 Dynamic Potential Distribution
Sagawa et al. (2019) consider robustness to arbitrary subpopulations, which is overly conservative and degenerates into training only on the highest-loss domain. To address this issue, Zhou et al. (2021) propose a more reasonable assumption by restricting $\mathcal{Q}$ in Equation 2 to an $f$-divergence ball (Csiszár, 1967) around a reference data ratio $\mathbf{p}_{R}$. This yields promising results, better ensuring domain balance (Jiao et al., 2022). Formally,
$\mathcal{Q}=\left\{\mathbf{q}:\chi^{2}\left(\mathbf{q},\mathbf{p}_{R}\right)\leq\rho\right\}.$ (5)
@@
To address this, we propose a method that combines the strengths of the aforementioned approaches. We still employ Equation 5 to constrain the distribution within a limited range, while gradually shifting the reference data ratio towards domains with higher losses to improve the model’s robustness to more challenging distributions. To ensure adequate training across all traversed potential distributions, we gradually update the reference ratio.
@@
$\mathbf{p}^{t+1}_{R}=\delta\cdot\mathbf{q}^{t}+(1-\delta)\cdot\mathbf{p}^{t}_{R}.$ (6)
@@
Overall process.
As illustrated in Figure 1, we treat each evaluation phase as an iteration. First, the evaluation loss is used to update the reference loss. Then, using the new reference loss along with the previous reference data ratio as input to the DRO process, we obtain a new data proportion for the next training step and update the reference data ratio accordingly.
@@ 4.1 Experimental Setup
To ensure comparability with Sheared Llama, we maintain consistency with most of its experimental settings. Further details of our experimental setup are provided in Appendix A.1.
@@
We allocate 0.4 billion tokens for pruning, utilizing the publicly available pruning dataset of Sheared Llama. We employ 50 billion tokens for continued pretraining. To enhance model quality, we use SlimPajama (Shen et al., 2023), a filtered version of RedPajama (Computer, 2023), and use its training split for continued pretraining.
Model configurations.
We use the Llama2-7B model (Touvron et al., 2023b) as the base model throughout our main experiments. We employ the same target architecture as Sheared Llama for structured pruning to ensure a fair comparison. We compare our results to strong open-source models of similar sizes, including Pythia-1.4B and 2.8B (Biderman et al., 2023) and Sheared Llama-1.3B and 2.7B (Xia et al., 2024). Additionally, we reproduce Sheared Llama, using the same data settings to control for other variables (ReSheared).
@@
To further explore the potential applications of the base model, we follow Sheared Llama by training with 10,000 instruction-response pairs sampled from the ShareGPT dataset and using another 1,000 instructions for evaluation. We follow Wang et al. (2024) and Xia et al. (2024) to employ LLMs, specifically GPT-4o, as an evaluator to compare the responses of the two models and report the win rates.
Method From To PPL $\downarrow$ Down. $\uparrow$
@@
Table 1: Perplexity (PPL) and downstream task performance (Down.) of pruned models. “Sheared Llama” is the open-source model, “ReSheared” is our reproduction of Sheared Llama, and “DRPruning” refers to the method proposed in this paper. “From” and “To” denote model size before and after pruning.
Figure 2: The curve of PPL changes during pruning from 7B for our reproduction of Sheared Llama and our method. Over the first 640 iterations (the vertical dash line), the model size is gradually reduced from 7B to the target size, which causes an initial increase in PPL.
@@
Llama2† Pythia† Sheared† ReSheared DRPrun. Pythia† Sheared† ReSheared DRPrun.
@@
Table 2: Our method (DRPrun.) outperforms publicly available models of comparable size on downstream tasks, as well as the Sheared Llama strategies. “Sheared” is the open-source model, while “ReSheared” is our reproduction. The number of shots is indicated in parentheses, with 0-shot used if unspecified. A model marked with †  denotes pretraining on different data. ReSheared and our method share the same experimental setup.
@@ 4.2 Main Results for Pruning
Our method surpasses the Sheared Llama during pruning in PPL and downstream tasks.
Our method significantly accelerates convergence.
We record the average perplexity (PPL) across different domains on the validation set. Table 1 shows that after 3.2k steps, PPL drops by 2.60% compared to the open-source model and by 5.59% in a fair comparison. Figure 2 confirms a more stable PPL as pruning proceeds, with further potential in later training stages, suggesting additional gains with extended training.
Our method better preserves the original model’s performance during pruning.  For a comprehensive analysis, we evaluate the model’s downstream task performance post-pruning. Across 15 tasks, our method consistently improves results, with an average gain of +1.53% over the open-source model and +1.27% in a fair comparison.
@@
4.3 Main Results for Continued Pretraining
We show that our method outperforms existing LLMs and other pruning approaches on both standard LM benchmarks and instruction tuning.
The LLMs recovered by our method are better foundation models after continued pretraining.
Table 2 presents the downstream performance of models with similar sizes, showing that our model surpasses most open-source models. Although we cannot replicate the 2.7B Sheared Llama results due to data differences, our model still outperforms most open-source LLMs, achieving +4.95% on average. Additionally, under consistent experimental conditions, our method outperforms our reproduction of Sheared Llama, achieving +1.78% on average. Moreover, we achieve significantly lower perplexity: 5.61 vs 5.97 for 1.3B, and 5.00 vs 5.27 for 2.7B, averaging a -5.60% reduction.
The effectiveness of our method is further demonstrated by supervised finetuning.
We compare the win rates of our instruction-tuned model with Sheared Llama. Figure 3 shows our model achieves a 55.4% win rate, surpassing both the open-source Sheared Llama and our reproduction. This highlights our model offers a stronger foundation for further use.
Figure 3: Win rate during instruction tuning. “Sheared Llama” refers to the open-source version, while “ReSheared” denotes our reproduced version. Our approach consistently outperforms Sheared Llama across all settings.
Figure 4: Effectiveness of our method compared to constant scheduling and naïve DRO during the 1.3B continued pretraining. The left figure shows PPL trends and the right shows average performance across 15 downstream tasks.
Figure 5: Variation of data ratio, reference data ratio, and reference loss during continued pretraining in the main experiment. The top three plots show results for pruning to 1.3B parameters, while the bottom three are for pruning to 2.7B parameters. Dashed lines in the data ratio plots represent the initial reference data ratio for each domain. Reference loss plots display the difference from the initial value, with the dashed line at $y=0$ indicating no change from the initial reference loss.
5 Analysis
5.1 Ablation Study
Our method significantly outperforms the DRO method.
As shown in Figure 4, our method consistently reduces PPL compared to the constant approach (-0.067 by the end of training), while the DRO method shows a slight increase (+0.011). In downstream tasks, our approach also outperforms both baselines, with performance improving steadily in the mid to late stages of training (-0.26 for DRO, +0.95 for ours). This underscores the sensitivity of DRO to hyperparameters and the necessity of dynamic adjustments.
@@ 5.1 Ablation Study
Our method dynamically identifies underperforming domains.

### v2

@@ Abstract
Large language models (LLMs) deliver impressive results but face challenges from increasing model sizes and computational costs. Structured pruning reduces model size and speeds up inference but often causes uneven degradation across domains, leading to biased performance. To address this, we propose DRPruning, a method that dynamically adjusts the data distribution during training to restore balanced performance across heterogeneous and multi-tasking data. Experiments in monolingual and multilingual settings show that DRPruning surpasses similarly sized models in both pruning and continued pretraining over perplexity, downstream tasks, and instruction tuning. Further analysis demonstrates the robustness of DRPruning towards various domains and distribution shifts. Furthermore, DRPruning can determine optimal reference losses and data ratios automatically, suggesting potential for broader applications. Code and scripts are available at https://github.com/hexuandeng/DRPruning.
@@
Hexuan Deng1    Wenxiang Jiao    Xuebo Liu1thanks:   Corresponding Author    Jing Li1    Min Zhang1    Zhaopeng Tu 1Institute of Computing and Intelligence, Harbin Institute of Technology, Shenzhen, China {hxuandeng,wenxiangjiaonju,tuzhaopeng}@gmail.com, {liuxuebo,li.jing,zhangmin2021}@hit.edu.cn
@@ 1 Introduction
Large language models (LLMs) have advanced rapidly, achieving impressive results across a wide range of tasks (Jiao et al., 2023b; Frieder et al., 2023; Deng et al., 2023; Bian et al., 2024; Rao et al., 2024). However, this progress has come with increasing model sizes, significantly raising computational costs for both training and inference, which impacts their accessibility. Structured pruning is a promising approach to reduce model size (Han et al., 2015; Wen et al., 2016), but it often causes uneven performance degradation across domains, leading to biased capabilities and unfair downstream task performance (Xia et al., 2024).
Given that LLMs inherently handle heterogeneous, multi-domain data, distribution robustness becomes essential. A commonly used approach is distributionally robust optimization (DRO; Oren et al., 2019; Sagawa et al., 2019), which aims to optimize worst-case performance across distributions. A reference loss is defined for each domain as a target. Domains with larger deviations from this reference loss are assigned higher weights, while not straying too far from a predefined reference data ratio. However, setting these hyperparameters is challenging, and suboptimal configurations often result in poor outcomes (Zhou et al., 2021).
To address this, we propose DRPruning, a distributionally robust pruning method that incorporates DRO to dynamically adjust the data distribution during training. Further, using scaling laws (Kaplan et al., 2020; Ghorbani et al., 2022), we predict the loss after training as the reference loss, where larger deviations indicate poorer performance, thereby promoting capability recovery in these areas. Additionally, we gradually increase the reference data ratio for domains with greater deviations, ensuring robustness across a wider range of distributions, particularly more challenging ones.
DRPruning is validated through experiments in monolingual and multilingual settings, which represent varying degrees of distributional shift. DRPruning outperforms other data scheduling methods in both pruning and continued pretraining, as measured by perplexity (-5.59%), downstream tasks (+1.52%), and instruction tuning (55.4% win rate). Particularly in multilingual settings, DRPruning achieves +2.95% in downstream tasks. To further assess domain-specific performance, we develop a sentence continuation benchmark using existing unlabeled data, demonstrating our improved domain-level capabilities (+17.9%).
@@
  • DRPruning tackles domain imbalance in structured pruning by introducing a distributionally robust pruning method, that dynamically adjusts data ratios during training to ensure robustness against distributional shifts.
  • We validate DRPruning through extensive experiments in monolingual and multilingual settings. Further analysis confirms its advantages in handling data heterogeneity and distribution shifts.
  • DRPruning offers refined reference losses and data ratios, which can be applied more broadly to enhance various model training processes and contribute to advancements for LLMs.
@@ 2.1 Structured Pruning
To prune the model to any target configuration, we adopt structured pruning based on Sheared Llama (Xia et al., 2024). For each granularity $i$, pruning masks $Z=\{\mathbf{z}^{i}\mid\mathbf{z}^{i}\in\mathbb{R}^{D_{i}}\}$ are learned to determine whether substructures are pruned or retained, where $z^{i}_{j}=0$ indicates pruning of the $j$-th substructure. Pruning is applied at various granularities, including transformer layers, hidden dimensions, attention heads, and FFN intermediate dimensions.
To parameterize the masks, the $\ell_{0}$ regularization method (Louizos et al., 2018) with hard concrete distributions is used to concentrate probability mass at 0 or 1. Lagrange multipliers are then used to ensure the pruned model meets the target configuration. Specifically, if exactly $t^{i}$ parameters must be retained for $\mathbf{z}^{i}$, the following constraint is imposed:
~$\tilde{\ell}^{i}=\lambda^{i}\left(\sum_{j}z^{i}_{j}-t^{i}\right)+\phi^{i}\left(\sum_{j}z^{i}_{j}-t^{i}\right)^{2}.$
The final training loss integrates these constraints with the language modeling loss of the pruned model, jointly optimizing the model parameters $\theta$ and pruning masks $\mathbf{z}$, with $\mathbf{z}$ typically uses a higher learning rate. After pruning, the highest-scoring components are retained.
@@ 2.2 Distributionally Robust Optimization
To mitigate uneven domain performance after pruning, we apply distributionally robust optimization (DRO; Oren et al., 2019; Sagawa et al., 2019) to improve the model’s robustness to distribution shifts. DRO seeks a model $\theta$ that performs well across a set of potential test distributions $\mathcal{Q}$ over $n$ domains. Formally:
@@
To solve the min-max optimization, the iterative best response algorithm (Fudenberg and Levine, 1998) is used. Each iteration consists of first performing the empirical risk minimization on the current data distribution $\mathbf{q}^{t}$, followed by updating the data distribution using worst-case weights based on the current parameters. Formally,
@@
3 Our Proposed DRPruning Method
To address the challenges of LLMs in handling heterogeneous and multi-tasking data, we propose DRPruning, illustrated in Figure 1. Each evaluation phase is treated as an iteration: the evaluation loss first updates the reference loss, which, along with the previous reference data ratio, serves as input to the DRO process. This yields a new data proportion for the next training step, and the reference data ratio is updated accordingly.
@@ 3.1 Distributionally Robust Pruning
We first introduce the overall procedure by applying naïve DRO to design an effective pruning and continued pretraining method. Specifically, we adopt common techniques (Ma et al., 2023; Li et al., 2024), first applying structured pruning to reduce model parameters, followed by continued pretraining to restore capabilities. Compared to training from scratch, this approach requires fewer unlabeled data to restore the model’s performance (Zhang et al., 2024b).
Integrate DRO into pruning and continued pretraining.
During training, we use DRO to dynamically adjust the data ratio to improve the model’s robustness and convergence speed. Specifically, to prevent overfitting, we compile a validation set and use the evaluation loss as the loss score. After each evaluation, we update the data ratio based on the evaluation loss using the DRO method, as in Eqn. 3, guiding training to focus more on underperforming domains.
@@
Next, we optimize the loss function $\ell$ (Section 3.2) and potential distributions $\mathcal{Q}$ (Section 3.3) to ensure robust training, as shown in Figure 1. In contrast, Sheared Llama employs a dynamic scheduling strategy that forces the model to strictly adhere to the relative loss magnitudes of larger LLMs, without placing any constraints on the potential distributions. This leads to suboptimal results, particularly in multilingual settings with significant distribution shifts.
@@ 3.2 Dynamic Loss Function
^$\hat{\ell}(P,T)=A\cdot\frac{1}{P^{\alpha}}\cdot\frac{1}{T^{\beta}}+E,$
@@ 3.3 Dynamic Potential Distribution
Sagawa et al. (2019) consider robustness to arbitrary subpopulations, which is overly conservative and degenerates into training only on the highest-loss domain. To address this issue, Zhou et al. (2021) propose a more reasonable assumption by restricting $\mathcal{Q}$ in Eqn. 2 to an $f$-divergence ball (Csiszár, 1967) around a reference data ratio $\mathbf{p}_{R}$. This yields promising results, better ensuring domain balance (Jiao et al., 2022). Formally,
$\mathcal{Q}=\left\{\mathbf{q}:\chi^{2}\left(\mathbf{q},\mathbf{p}_{R}\right)\leq\rho\right\}.$
@@
To address this, we propose a method that combines the strengths of the aforementioned approaches. We still employ Eqn. 5 to constrain the distribution within a limited range, while gradually shifting the reference data ratio towards domains with higher losses to improve the model’s robustness to more challenging distributions. To ensure adequate training across all traversed potential distributions, we gradually update the reference ratio.
@@
$\mathbf{p}^{t+1}_{R}=\delta\cdot\mathbf{q}^{t}+(1-\delta)\cdot\mathbf{p}^{t}_{R}.$
@@ 4.1 Experimental Setup
Model.
Llama2-7B model (Touvron et al., 2023b) is used as the base model. We employ the same target architecture as Sheared Llama for structured pruning to ensure a fair comparison. We compare our method, i.e., DRPruning, to strong open-source models of similar sizes, including Pythia-1.4B and 2.8B (Biderman et al., 2023) and Sheared Llama-1.3B and 2.7B. Additionally, we reproduce Sheared Llama, using the same data settings to control for other variables (ReSheared). Further details are provided in Appendix A.1.
@@
To ensure comparability with Sheared Llama, we align most of our settings with its approach. However, due to insufficient documentation of its data filtering method, we are unable to replicate the results under the 2.7B setting. Therefore, our comparison primarily focuses on ReSheared and DRPruning, using a similar data setting for our reproduction. We allocate 0.4 billion tokens for pruning, utilizing the publicly available pruning dataset of Sheared Llama. We employ 50 billion tokens for continued pretraining, and use SlimPajama (Shen et al., 2023), a filtered version of RedPajama (Computer, 2023), and use its training split for continued pretraining.
@@
To further explore the potential applications of the base model, we follow Sheared Llama by training with 10k instruction-response pairs sampled from the ShareGPT dataset and using another 1k instructions for evaluation. We follow Wang et al. (2024) and Sheared Llama to employ LLMs, specifically GPT-4o, as an evaluator to compare the responses of the two models and report the win rates.
Method From To PPL $\downarrow$ Task $\uparrow$
@@
Table 1: Perplexity (PPL) and downstream task performance (Task) of pruned models. The “Task” performance represents the macro-average across 15 tasks, as detailed in Section 4.1. “From” and “To” indicate the model size before and after pruning, respectively.
Figure 2: The curve of PPL changes during pruning from 7B. Over the first 640 iterations (the vertical dash line), the model size is gradually reduced from 7B to the target size, which causes an initial increase in PPL.
@@
Llama2 Pythia Sheared ReSheared DRPrun. Pythia Sheared ReSheared DRPrun.
@@
Table 2: Performances of different models across 15 downstream tasks. “Sheared” refers to Sheared Llama. “ReSheared” is our reproduction of Sheared Llama. “DRPrun.” refers to our method. The number of shots is indicated in parentheses, with 0-shot used when unspecified. A model marked with †  indicates training on different data. Bold and underlined represent the best and second-best results, respectively, for each model size.
@@ 4.2 Main Results for Pruning
Our method surpasses ReSheared during pruning in PPL and downstream tasks.
DRPruning promotes convergence by increasing the weight of underperformed domains.
To demonstrate this, we record the average PPL across different domains on the validation set. Table 1 shows that our method achieves lower PPL. Figure 2 confirms faster convergence as pruning proceeds, with further potential in later training stages, suggesting additional gains with extended training.
Our method better preserves the original model’s performance during pruning.
For a comprehensive analysis, we evaluate the model’s downstream task performance post-pruning. Across 15 tasks, our method significantly improves performance on average, with +1.53% over the open-source model and +1.27% in a fair comparison.
@@
4.3 Main Results for Continued Pretraining
DRPruning outperforms ReSheared on average across LM benchmarks and instruction tuning.
LLMs recovered by our method are better foundation models after continued pretraining.
Table 2 presents the downstream performance of models with similar sizes, showing that our model surpasses most open-source models. We are unable to replicate the results of the 2.7B Sheared Llama due to differences in data, where our approach yields worse performance. Nevertheless, we outperform other open-source LLMs. On average, we achieve an improvement of +4.95% comparing to open-source LLMs. Additionally, under consistent experimental conditions, our method outperforms ReSheared, achieving +1.78% on average. Its statistical significance is confirmed by t-test in Appendix B.3. Moreover, we achieve significantly lower perplexity: 5.61 vs 5.97 for 1.3B, and 5.00 vs 5.27 for 2.7B, averaging a -5.60% reduction.
Figure 3: Win rate during instruction tuning. DRPruning outperforms Sheared Llama and ReSheared.
Figure 4: Effectiveness of our method compared to constant scheduling and naïve DRO during the 1.3B continued pretraining. Left figure: PPL trends; Right: average performance across 15 downstream tasks.
The effectiveness of DRPruning is further demonstrated by instruction tuning.
We compare the win rates of our instruction-tuned model with Sheared Llama and ReSheared. Figure 3 shows our model achieves a 55.4% win rate, surpassing both the open-source Sheared Llama and ReSheared. This highlights that DRPruning offers a stronger foundation for further use.
Additionally, DRPruning introduces no extra GPU computation, with the only overhead stemming from data ratio calculation, contributing to less than 1.5% of the training time. Furthermore, we implemented parallel data ratio computation and demonstrated that it does not impact performance, as detailed in Appendix B.4.
5 Analysis
5.1 Ablation Study
DRPruning significantly outperforms DRO method.
As shown in Figure 4, our method consistently reduces PPL compared to the constant and DRO approach. In downstream tasks, DRPruning also outperforms both baselines, with performance improving steadily in the mid to late stages of training (-0.26 for DRO, +0.95 for ours). This underscores the sensitivity of DRO to hyperparameters and the necessity of dynamic adjustments.
Our strategy dynamically identifies underperforming domains.